OpenStudy (anonymous):

Solve. 3 3 x ----- = ----- + ----- x-3 2 x-2

OpenStudy (anonymous):

first Simplify 3/2 + x/x-2 (3*-1+x)/(2*-1+x)=x-3/x-2 (x-3/x-2)+x/(x-2)=(2x-3)/x-2 3/(x-3)=(2x-3)/x-2 3*(x-2)=(x-3)*(2x-3) When you complete the simplification you get x=3/2

OpenStudy (anonymous):

x=3/2